OpenStudy (anonymous):

help i got 20 mins till its dew What is the ratio 14/35 expressed in simplest form?

OpenStudy (alexandervonhumboldt2):

hint: 14=2*7 35=7*5

OpenStudy (anonymous):

its a typeing one so no mulit choise

OpenStudy (alexandervonhumboldt2):

\[\frac{ 14 }{ 35 }=\frac{ \cancel 7*2 }{ \cancel 7*5 }=\]

OpenStudy (anonymous):

2/5?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so 2/5 thx